Required information E9-2 Computing and Recording a Basket Purchase and Straight-Line Depreciation [LO 9.2, LO 93] [The following information applies to the questions displayed below Bridge City Consulting bought a building and the land on which it is located for $182,000 cash. The land is estimated to represent 70 percent of the purchase price. The company paid $22,000 for building renovations before it was ready for use. E9-2 Part2 Required: 2. Prepare the journal entry to record all expenditures. Assume that all transactions were for cash and they occurred at the start of the year. (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field.) View transaction list Journal entry worksheet Record all expenditures for the land and buildings assuming all transactions were paid for with cash and occurred at the start of the year Prev 1 2 of 8Next >
0 Required information Required 2. Prepare the journal entry to record all expenditures. Assume that all transactions were for cash and they occurred at the start of t year. (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field.) View transaction list Journal entry worksheet Record all expenditures for the land and buildings assuming all transactions were paid for with cash and occurred at the start of the year Note: Enter debits before credits. Transaction Debit | Credit Record entry Clear entry View general journal (Prev 12 of 8lill: Next >
Required information E9-2 Computing and Recording a Basket Purchase and Straight-Line Depreciation [LO 9-2, LO 9-3 The following information applies to the questions displayed below. Bridge City Consulting bought a building and the land on which it is located for $182,000 cash. The land is estimated to represent 70 percent of the purchase price. The company paid $22,000 for building renovations before it was ready for use E9-2 Part 3 and 4 3. Compute straight-line depreciation on the building at the end of one year, assuming an estimated 12-year useful life and a $4,600 estimated residual value. (Do not round intermediate calculations.) 4. What should be the book value of (e) the land and (b) the building at the end of year 2? 4(a) Land 4(b) Building
Answer #1

2) Journal entry

Date account and explanation debit credit
Land (182000*70%) 127400
Building (182000*30%+22000) 76600
Cash 204000
(To record purchase assets)

3) Straight line depreciation = (76600-4600)/12 = 6000 per year

4a) Book value of land = 127400

Book value of building = 76600-6000-6000 = 64600